Circular No. 5794

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA AQUILAE 1993 D. Balam reports the following precise position obtained with the 0.5-m

reflector (+ CCD) at the University of Victoria’s Climenhaga Observatory (mean error 0′′.2): α = 19^{h}10^{m}34^{s}.69, δ = +1^{o}29′14′′.0 (equinox 1950.0). Photographic magnitude estimates from Technical Pan film (+ orange filter) by W. Liller, Viña del Mar, Chile: Apr. 26.244 UT, [11.5; May 13.351, 8.7 (independent discovery); 17.246, 8.2.

The following Johnson photoelectric data have been reported: May 17.279 UT, V = 7.90 ± 0.08, B–V = +0.89 ± 0.05 (D. Kaiser, Columbus, IN; comparison star HD 178065); 17.306, V = 7.90 ± 0.006, B–V = +0.91 ± 0.008 (Kaiser); May 18.055, V = 8.058 ± 0.004, U–B = +0.03, B–V = +0.83, V–R = +0.97, V–I = +1.76 (M. Mikołajewski and B. Wikierski, N. Copernicus University, Toruń; comparison star HR 7313); 18.096, V = 8.22 ± 0.03, B–V = +0.41 ± 0.05 (H. Mikuž, Ljubljana, Slovenia; comparison star HR 7313, V = 6.19, B–V = +0.02).

PSR 1259–63 S. Johnston and R. Manchester, Australia Telescope National Facility;

A. Lyne, University of Manchester; and N. D’Amico, University of Palermo, report: “We have improved parameters for the binary system containing the pulsar PSR 1259–63 and the Be star SS 2883 (Johnston et al. 1992, Ap.J. 387, L37), obtained from a phase-connected solution of 1200 days of radio timing data from the Parkes radio telescope. The solution yields period 47.762 053 92 ms and period derivative 2.280 × 10^{−15} at epoch 1990 June 11.0 UT. Binary orbital parameters are a(sin i) = 1296.00 light seconds, P = 1236.81 days, e = 0.8698, ω = 138^{o}.655, epoch of periastron 1990 Aug. 21.36 UT. The mass function leads to a minimum companion mass of 3.2 M_{⊙}. Based on optical observations, we favor solutions with a companion mass of ∼ 10 M_{⊙}, implying that the inclination angle of the orbit to the plane of the sky is ∼ 36^{o}. The next periastron approach is 1994 Jan. 9. During periastron, the pulsar may disrupt the circumstellar disk of the Be star, resulting in possible x-ray and γ-ray emission and changes in the optical brightness and spectral-line profiles of SS 2883. Monitoring of this system is encouraged around the epoch of next periastron.”

1993 May 18 Daniel W. E. Green